Find the general solution using D-operator

(D + 4)^2x = sihn4t


1
Expert's answer
2022-04-16T04:13:03-0400

Solution

For homogeneous equation (D + 4)2x =  0 the characteristic equation is

(λ + 4)2=0 =>  λ1,2 = -4

So the solution of homogeneous equation is x0(t) = C1e-4t+C2te-4t, where C1, C2 are arbitrary constants.

For given nonhomogeneous equation

(D + 4)2x = sinh4t  or  (D + 4)2x = 0.5e4t – 0.5e-4t     

partial solution may be found in the form

x1(t)= Ae4t + Bt2e-4t        

Substitution this into equation gives

(D + 4)x1 =8Ae4t + 2Bte-4t  =>  

(D + 4)2x1 = (D + 4)(8Ae4t + 2Bte-4t ) = 64Ae4t + 2Be-4t  =>         

64Ae4t = 0.5, 2Be-4t = -0.5  => A = 1/128, B = -1/4

So general solution of given equation is

x(t) = x0(t) + x1(t) = C1e-4t+C2te-4t + e4t/128 - t2e-4t/4